### v1

# Optimizing Sequencing Coverage Depth in DNA Storage: Insights From DNA Storage Data

Abstract
DNA data storage is now being considered as a new archival storage method for its durability and high information density, but still facing some challenges like high costs and low throughput. By reducing sequencing sample size for decoding digital data, minimizing DNA coverage depth helps lower both costs and system latency. Previous studies have mainly focused on minimizing coverage depth in uniform distribution channels under theoretical assumptions. In contrast, our work uses real DNA storage experimental data to extend this problem to log-normal distribution channels, a conclusion derived from our PCR and sequencing data analysis. In this framework, we investigate both noiseless and noisy channels. We first demonstrate a detailed negative correlation between linear coding redundancy and the expected minimum sequencing coverage depth. Moreover, we observe that the probability of successfully decoding all data in a single sequencing run increases and then decreases as coding redundancy rises, when the sample size is optimized for complete decoding. Then we extend the lower bounds of DNA coverage depth from uniform to log-normal noisy channels. The findings of this study provide valuable insights for the efficient execution of DNA storage experiments.
Index Terms:
DNA data storage, sequencing coverage depth, log-normal distribution channel.
I Introduction
With the rapid growth of global data and advancements in information technology, traditional storage media such as HDDs and SSDs may no longer meet future data storage demands [1, 2]. DNA storage, with its high density and durability, has emerged as a promising solution to address this challenge. However, current DNA storage technologies face significant obstacles, including system latency due to low throughput and high costs. Minimizing sequencing coverage depth in DNA storage systems could provide valuable insights for reducing latency and associated costs.
Several studies have addressed the issue of sequencing coverage depth in non-random access scenarios, focusing on minimizing sequencing coverage depth in channels follow uniform distribution by adjusting outer error-correcting codes. These approaches are often based on classical problems such as the coupon collector’s problem, the urn problem, and others. [3] first introduced the concept of sequencing coverage depth, investigating the minimum coverage depth required for successful message decoding under non-random access settings. They also explored the singleton coverage depth problem and proposed a new coding scheme to reduce the sample size needed to retrieve a single strand. [4] extended the work of [3] to more practical application scenarios, particularly focusing on retrieving $k$ (where $k=1$ or $2$) out of $n$ files. This study also examined the structural properties of various coding schemes and their impact on random access expectations and probability distributions. [5] provided a novel perspective on random access efficiency by exploring coding matrix design and performance optimization. Their work introduced the concept of recovery balanced codes and corresponding conditions, offering theoretical guidance for designing more efficient DNA storage systems. [6] proposed the first model to analyze sequencing coverage depth in combinatorial coding, calculating the probability of error-free reconstruction and investigating theoretical bounds to evaluate decoding probabilities. These studies collectively address the problem of minimizing sequencing coverage depth in uniformly distributed channels, offering various theoretical models, coding schemes, and analytical methods. Their contributions provide a foundation for optimizing sequencing efficiency and cost-effectiveness in DNA storage systems. However, the processes of DNA synthesis and amplification exhibit a degree of randomness, which results in a non-uniform channel probability distribution. Therefore, in this study, we investigate the problem of minimizing sequencing coverage depth in a real-world channel based on PCR and sequencing data, under the non-random access setting.
This paper is organized as follows: Section II provides an overview of related work and offers a detailed description of the problem addressed in this study. Section III analyzes PCR and sequencing data to establish the foundation for research on minimizing sequencing coverage depth in channels where the probability distribution follows a log-normal distribution. Section IV investigates the expected value of minimum sequencing coverage depth in the noiseless channel and two lower bounds in the noisy channel, under the non-random access setting.
II Problem Statement, Related Work
II-A Problem Statement
The problem we study is built on the following DNA storage model (see Fig. 1).
Figure 1: Framework of this paper.
For alphabet $\varSigma=\{A,C,G,T\}$, each DNA strand or sequence of length $l$ can be represented as a vector of length $l$ over the alphabet $\varSigma$. The set of all such vectors of length $l$ over $\varSigma$ is denoted by $\varSigma^{l}$. Let $\varSigma^{*}\triangleq\bigcup_{l=0}^{\infty}\varSigma^{l}$.
Write and encode. Assume that the digital data to be stored is first segmented and translated into $m$ information strands of length $l$, denoted as $X=(x_{1},x_{2},\dots,x_{m})\in(\varSigma^{l})^{m}$. After encoding with error-correcting code $\mathcal{C}$, the $m$ information strands are transformed into $n$ designed strands of length $l$, represented as $D=(d_{1},d_{2},\dots,d_{n})\in(\varSigma^{l})^{n}$. Next, each designed strand in $D$ is synthesized into $c$ copies. Due to the inherent randomness in the synthesis process, the actual copy numbers of each designed strand $c_{i}$ is not the same [12], where $i\in[1,n]$, and each synthesis strand is a noisy version of its corresponding designed strand. Let $M=\sum_{i=1}^{n}c_{i}$, then, all of $M$ synthetic strands will be stored unordered in a container until reading them out.
Read and decode. In this step, we will extract a specific volume or weight of DNA dry powder in the container and performing sequencing using next-generation sequencer (NGS). Let the multiple set $S_{K}=\{\{s_{1},s_{2},\dots,s_{K}\}\}$ represents the $K$ reads obtained from sequencing, where each $s_{i}\in\varSigma^{*}$ is a noisy copy of a particular designed strand $d_{j}$. During the sequencing process, due to differences in sequencing technologies, the reads in $S_{K}$ can either be obtained sequentially, as in Nanopore sequencing, or all at once, as in Illumina sequencing. However, the research conclusions drawn from both approaches do not show significant differences [7]. In this study, a sequential reading method is employed. The number of noisy copies of the $i$-th designed strand $d_{i}$ in $S_{K}$ depends on the channel probability distribution $\boldsymbol{p}_{t}=\left(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)}\right)$, where $p_{i}^{(t)}$ represents the probability of obtaining a copy of $d_{i}$ in each round, i.e., the proportion of $d_{i}$ in the population after certain cycles of PCR amplification. During the decoding process, we first apply a clustering algorithm to partition the reads in $S_{K}$ into clusters based on their corresponding designed strands. Specifically, the $i$-th cluster contains all noisy copies of the $i$-th designed strand $d_{i}$ in $S_{K}$, denoted as $s_{J}=\{s_{j}|s_{j}\ \text{is a noisy copy of}\ d_{j}\}$. Based on the assumptions and analysis outlined above, the probability of successfully retrieving both the information strand set $X$ and the designed strand set $D$ primarily depends on the error-correcting code $\mathcal{C}$ and the retrieval algorithm [7].
  1. 1.
    Explanation of error-correcting code $\mathcal{C}$. The construction of the error-correcting code determines how many designed strands need to be retrieved to decode all the digital data. In our work, for given $m$ information strands, we assume the use of a $[n,m]$ MDS code $\mathcal{C}$ with redundancy $R=\frac{n}{m}$, meaning that during the decoding process, retrieving only $m$ out of $n$ designed strands is sufficient to decode the entire dataset.
  2. 2.
    Explanation of retrieval algorithm. The retrieval algorithm determines how many noisy copies of a designed strand need to be retrieved in order to successfully decode that strand. In our work, we model the retrieval algorithm using a positive integer $a\geq 1$, assuming that at least $a$ noisy copies of $d_{i}$ need be retrieved for $d_{i}$ to be successfully decoded.
Notice that due to the inherent randomness in the synthesis, amplification, and sequencing processes of DNA storage [13], it is uncertain whether all the original information can be fully recovered from the $K$ reads obtained through sequencing. Thus, we focus on how to ensure the 100% successful decoding of the original data with high probability, using as small sequencing sample size as possible in the real channel.
To this end, we model the sample size as a function of the channel probability distribution $\boldsymbol{p}_{t}=\left(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)}\right)$, the [n,m] MDS code $\mathcal{C}$, and the retrieval algorithm $a$, where $p_{i}$ represents the probability that a single read is from the $i$-th designed strand after $t$ cycles of PCR amplification, and $a$ is a positive integer. We define a new notation $K\triangleq I_{a}^{\boldsymbol{p}_{t}}(n,m)$, which means the sample size $K$ to decode the full digital data under the condition that after $t$ cycles of PCR, the retrieval algorithm needs to retrieve at least $a$ noisy copies of each of $m$ out of $n$ designed strands. For a uniformly distributed channel, we will omit $\boldsymbol{p}_{t}$, and denote it as $I_{a}(n,m)$.
The main problems we investigate in this paper are defined below.
Problem 1. (Simulation of DNA storage channel distribution.) Given values of $n\geq m\geq 1$, $t\geq 1$, the synthesis amount of each designed strand $c_{1},c_{2},\dots,c_{n}\stackrel{{\scriptstyle i.i.d}}{{\sim}}p(c)$, and the amplification efficiency of each designed strand $r_{1},r_{2},\dots,r_{n}$, we focus on the following questions:
  1. 1.
    The probability distribution of $\nu_{i}^{(t)}$, i.e., copy numbers of the $i$-th designed strand after $t$ cycles of PCR amplification.
  2. 2.
    The probability distribution of $p_{i}^{(t)}=\frac{\nu_{i}^{(t)}}{\sum_{j=1}^{n}\nu_{j}^{(t)}}$, i.e., the proportion of the $i$-th designed strand in the population after $t$ cycles of PCR amplification.
Problem 2. (Coverage depth problem in the real channel.) Given values of $n\geq m\geq 1$, $a\geq 1$, we focus on the following questions:
  1. 1.
    The expectation value $\mathbb{E}[I_{a}(n,m)]$.
  2. 2.
    Given $t\geq 1$, the expectation value $\mathbb{E}[I_{1}^{\boldsymbol{p}_{t}}(n,m)]$.
  3. 3.
    Given $t\geq 1$, the lower bounds of $I_{a}^{\boldsymbol{p}_{t}}(n,m)$.
II-B Previous Work
By directly mapping the coupon collector’s problem, urn and dixie cup problem to the sequencing coverage depth problem, we obtain from [8, 9, 10, 11] that,
$\mathbb{E}[I_{1}(n,m=n)]=n\log n+\gamma n+\mathcal{O}(1),$
where $\gamma\sim 0.577$ is the Euler–Mascheroni constant.
$\mathbb{E}[I_{1}(n,m)]=n(H_{n}-H_{n-m}),$
where $H_{n}$ is the $n$-th harmonic number.
$\mathbb{E}[I_{a}(n=m,m)]$ $=m\log m+m(a-1)\log\log m$
$+mC_{a}+o(m),$
where $C_{a}$ is a constant that corresponding to $a$.
$\mathbb{E}[I_{a}^{\boldsymbol{p}}(n,m)]=\sum_{q=0}^{m-1}\int_{0}^{\infty}[v^{q}]\prod_{i=1}^{n}(e_{t-1}(p_{i}r)+v(e^{p_{i}r}-e_{t-1}(p_{i}r)))e^{-r}dr,$ (1)
where $\boldsymbol{p}$ represents any general discrete distribution, $[v^{q}]Q(v)$ represents the coefficients of the $q$-th terms of the polynomial $Q(v)$, and $e_{t}(x)=\sum_{i=1}^{t}\frac{x^{i}}{i!}$.
Based on the models above, previous work has been carried out mainly in channels following uniform distribution. For example, [3] proved that for any $\epsilon>0$,
$\log\left(\frac{1}{1-R}\right)+f_{c}(n,R)\leq\mathbb{E}\left[\frac{I_{a}(n,m)}{n}\right]\leq\left(\log\left(\frac{1}{1-R}\right)+a\log\log n+2\log(a+1)\right)\cdot(1+2\epsilon),$
where $f_{c}(n,R)=\mathcal{O}(\frac{1}{n^{2}})$.
Obviously, (1) is not a closed form expression and is not easy to calculate. Thus, we will build upon the MDS coverage depth problem in [3] that addresses the minimum coverage depth in uniform distribution channels under non-random access setting to derive the expectation value of $I_{a}^{\boldsymbol{p}_{t}}(n,m)$ in a noiseless channel, as well as its lower bounds in a noisy channel in Section IV.
III Simulation of DNA Storage Channel Probability Distribution
According to the definition of the channel probability distribution in Section II, it is closely related to the proportion of each strand in the population after PCR amplification. In this section, by analyzing PCR and sequencing data, we find that the real channel probability distribution follows a log-normal distribution. Based on this, we propose a theoretical model for the simulation of the real channel probability distribution, which serves as the foundation for the subsequent research on minimizing sequencing coverage depth.
III-A Analysis of Channel Probability Distribution Based on PCR and Sequencing Data
We first analyze the real channel probability distribution based on the PCR and sequencing data. We have an oligo pool of 11,520 distinct oligos(i.e., $n=11520$), and PCR amplification was performed using these oligos as template strands. After 10 cycles of amplification, sequencing 4,970,786 reads to generate Dataset PCR10; after 30 cycles, sequencing 11,001,029 reads to generate Dataset PCR30; and after 60 cycles, sequencing 11,180,177 reads to generate Dataset PCR60. The relationship between different PCR cycles and the copy numbers of different oligos at that cycle is shown in Fig. 4(a), 4(b) and 4(c).
Figure 2: Visualization of different cycles of PCR data.
Our primary interest lies in the relationship between the number of PCR cycles and the proportion of each strand in the population. Given that the sequencing data exhibit a clear skewed distribution, we assume that the PCR data follow a log-normal distribution [13]. Then we performed a logarithmic transformation on the normalized data before fitting a normal distribution, and Fig. 4(d), 4(e) and 4(f) show the results of a good fit on all three datasets, which means after PCR amplification, the channel actually follows a log-normal distribution.
Assuming $X$ is a random variable that follows a normal distribution with parameters $\mu$ and $\sigma^{2}$ as its expectation and variance after logarithmic transformation, we denote it as $X\sim\mathcal{LN}(\mu,\sigma^{2})$. After fitting the PCR data to a log-normal distribution, the sample distributions of the proportion of the $i$-th strand in the population after different PCR cycles are presented in Table I.
Since all three datasets are samples from the overall population generated by PCR amplification for 10, 30, and 60 cycles, we will next perform maximum likelihood estimation (MLE) to estimate the parameters of the population. Let $x_{1},x_{2},\dots,x_{n}$ be a simple random sample from the population $X\sim\mathcal{LN}(\mu,\sigma^{2})$, the MLE for $\mu$ and $\sigma^{2}$ are denoted as $\hat{\mu}_{MLE}$ and $\hat{\sigma}^{2}_{MLE}$ respectively. Then, from [14], we have
$\hat{\mu}_{MLE}=\frac{1}{n}\sum_{i=1}^{n}\ln{(x_{i})},$ (2)
and
$\hat{\sigma}^{2}_{MLE}=\frac{1}{n}\sum_{i=1}^{n}(\ln{(x_{i})}-\hat{\mu}_{MLE})^{2}.$ (3)
Based on the calculations from (2) and (3), the population distributions of the proportions of different strands after 10, 30, and 60 PCR cycles are presented in Table I.
TABLE I: The distribution of different cycles of PCR data.
Cycles Sample distribution Population distribution
PCR10 $\mathcal{LN}(-9.71,0.86^{2})$ $\mathcal{LN}(-9.72,0.74^{2})$
PCR30 $\mathcal{LN}(-9.91,0.98^{2})$ $\mathcal{LN}(-9.86,0.96^{2})$
PCR60 $\mathcal{LN}(-10.38,1.11^{2})$ $\mathcal{LN}(-10.25,1.38^{2})$
III-B Real Channel Probability Distribution Model
In the previous subsection, we analyzed PCR experimental data to obtain the real channel distributions after 10, 30, and 60 cycles of PCR amplification. However, in practical experiments, additional sequencing of PCR data to obtain the parameter information of the probability distribution is typically not performed. This necessitates the simulation of the channel probability distribution under the condition that only the synthetic amount $c_{i}$ and amplification efficiency $r_{i}$ of each designed strand are known.
Therefore, for given $c_{i}$ and $r_{i}$, where $i\in[1,n]$, we model the channel probability distribution as a function of PCR cycles. First, it is evident that after $t$ cycles of PCR amplification, the expected copy number of the $i$-th strand is
$\mathbb{E}\left[\nu_{i}^{(t)}\right]=c_{i}(1+r_{i})^{t}.$
Thus, we can derive the proportion of the $i$-th strand in the population after $t$ cycles of PCR amplification as
$p_{i}^{(t)}=\frac{c_{i}(1+r_{i})^{t}}{\sum_{j=1}^{n}c_{j}(1+r_{j})^{t}},$
then the expectation and variance of $p_{i}^{(t)}$ can be calculated, denoted by $\mathbb{E}\left[p_{i}^{(t)}\right]$ and $Var\left[p_{i}^{(t)}\right]$.
According to [14], we can calculate the parameters $\mu^{(t)}$ and $\sigma^{(t)}$ of log-normal distribution is that,
$\mu^{(t)}=\log{\left(\mathbb{E}\left[p_{i}^{(t)}\right]\right)}-\frac{1}{2}\log{\left(1+\frac{Var\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)},$ (4)
and
$\sigma^{(t)}=\sqrt{\ln{\left(1+\frac{Var\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)}}.$ (5)
Then from the conclusion we obtained in the previous subsection, we have
$p_{i}^{(t)}\sim\mathcal{LN}\left(\log{\left(\mathbb{E}\left[p_{i}^{(t)}\right]\right)}-\frac{1}{2}\log{\left(1+\frac{Var\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)},\ln{\left(1+\frac{Var\left[p_{i}^{(t)}\right]}{\mathbb{E}^{2}\left[p_{i}^{(t)}\right]}\right)}\right).$
Remark. For the simulation of theorems in the following section, we use the population distribution of PCR amplification and sequencing data from those three different cycles.
IV Sequencing Coverage Depth Problem in the Log-normal Distribution Channel
Under the condition of non-random access, the main goal of our study is to find the minimum sequencing sample size $K$ that ensures the successful decoding of all digital data in the real channel that following the log-normal distribution after PCR. In this section, we first prove the expected coverage depth in log-normally distributed noiseless channels, and propose the problem of decoding all data successfully in a single sequencing experiment using the expected coverage depth as the sample size. Then we prove the theoretical lower bounds of expected coverage depth in the real noisy channels. For given $m$ information strands, we denote $\alpha\triangleq\frac{K}{m}$ as the sequencing coverage depth.
IV-A Sequencing Coverage Depth in the Noiseless Channel
In practical experiments, in addition to focusing on the expected sequencing coverage depth required to decode all the digital information (i.e., $\mathbb{E}[\alpha]$), we are also concerned with the probability of successfully decoding all data in a single sequencing run (i.e., $Var[\alpha]$). Let $N$ represents the number of different designed strands when sequencing fixed $K$ reads.
Remark. In a noiseless channel, we assume that only one copy of each designed strand is sufficient for successful retrieval of that strand, i.e., $a=1$.
Theorem 1. For given any channel probability distribution $\boldsymbol{p}_{t}=(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)})$ and any $K\geq n\geq m\geq 1$, it holds that
$N\sim\mathcal{N}\Bigg{(}\sum_{i=1}^{n}\left(1-e^{-K\cdot p_{i}^{(t)}}\right),\sum_{i=1}^{n}e^{-K\cdot p_{i}^{(t)}}\left(1-e^{-K\cdot p_{i}^{(t)}}\right)-K\left(\sum_{i=1}^{n}p_{i}^{(t)}e^{-K\cdot p_{i}^{(t)}}\right)^{2}\Bigg{)}.$
Proof. This problem is equivalent to determining the number of urns containing at least one ball after throwing $K$ balls into $n$ labeled urns (one ball per round). Notice that the probability of an identical ball falling into each urn is not equal, and they are given by $\boldsymbol{p}_{t}=\left(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)}\right)$, where $p_{i}^{(t)}$ represents the probability of the ball falls into the $i$-th urn in each round.
The value of $\mathbb{E}[N]$:
${X_{i}^{(K)}}=\begin{cases}1,&{\text{the i-th urn contains at least one ball after $K$ rounds}},\\ 0,&{\text{the i-th urn remains empty after $K$ rounds}.}\end{cases}$
thus
$\mathbb{E}[X_{i}]=1-\left(1-p_{i}^{(t)}\right)^{K}\approx 1-e^{K\cdot p_{i}^{(t)}},$
when $p_{i}^{(t)}$ is sufficiently small. Suppose there are $N$ urns containing at least one ball after $K$ rounds, then given $K$,
$\mathbb{E}[N]=\mathbb{E}\left[\sum_{i=1}^{n}X_{i}\right]=\sum_{i=1}^{n}\mathbb{E}[X_{i}]=\sum_{i=1}^{n}\left(1-e^{K\cdot p_{i}^{(t)}}\right).$
The value of $Var[N]$: Inspired by [15], we first investigate the number of urns that remain empty after $K$ balls thrown into $n$ urns. Let $N_{0}$ represents the number of urns that remain empty after $K$ balls thrown into $n$ urns. According to [15],
$Var[N_{0}]=\sum_{i=1}^{n}\left(e^{-Kp_{i}^{(t)}}-e^{-2Kp_{i}^{(t)}}\right)-K\left(\sum_{i=1}^{n}p_{i}^{(t)}e^{-Kp_{i}^{(t)}}\right)^{2}.$
Thus
$Var[N]=Var[n-N_{0}]=Var[N_{0}]=\sum_{i=1}^{n}\left(e^{-Kp_{i}^{(t)}}-e^{-2Kp_{i}^{(t)}}\right)-K\left(\sum_{i=1}^{n}p_{i}^{(t)}e^{-Kp_{i}^{(t)}}\right)^{2}.$
By the central limit theorem,
$N\sim\mathcal{N}\left(\sum_{i=1}^{n}\left(1-e^{K\cdot p_{i}^{(t)}}\right),\sum_{i=1}^{n}\left(e^{-Kp_{i}^{(t)}}-e^{-2Kp_{i}^{(t)}}\right)-K\left(\sum_{i=1}^{n}p_{i}^{(t)}e^{-Kp_{i}^{(t)}}\right)^{2}\right).$
[…]
On the basis of Theorem 1, let $\boldsymbol{p}_{t}=\left(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)}\right)=\left(\frac{1}{n},\frac{1}{n},\dots,\frac{1}{n}\right)$, the following result is straightforward.
Theorem 2. For a uniform distribution channel, and any $K\geq n\geq m\geq 1$, it holds that
$N\sim\left(m\left(1-e^{-\alpha}\right),m(e^{-\alpha}-e^{-2\alpha}-\alpha e^{-2\alpha})\right).$
In summary, for given $m$ information strands, when sequencing fixed $K$ reads, the expected ratio of successfully decoded designed strands is $\frac{N}{m}$. That is, if $m$ information strands are encoded with $\left[\frac{m^{2}}{N},m\right]$ MDS code, then theoretically speaking, sequencing $K$ reads in a noiseless channel recovers the full information.
Applying the conclusions of Theorem 1 and Theorem 2, we present in Fig. 2 a comparison of the minimum sequencing coverage depth under real channels and a uniform distribution channel. Fig. 3 visualize the variance derivative function to intuitively show the probability of successfully decoding all the digital data in a single sequencing experiment, where $f(K)\triangleq e^{-K\mathbb{E}\left[p_{i}^{(t)}\right]}-e^{-2K\mathbb{E}\left[p_{i}^{(t)}\right]}-\frac{n}{K}\left(K\mathbb{E}\left[p_{i}^{(t)}\right]\right)^{2}e^{-2K\mathbb{E}\left[p_{i}^{(t)}\right]}$.
Based on the applications of the Theorem 1 and Theorem 2, we obtain the following conclusions:
  1. 1.
    When encoding $m$ information strands with identical redundancy, the minimum sequencing coverage depth required to decode all data in a uniform distribution channel is considerably smaller than in a real channel. Therefore, further investigation of the minimum sequencing coverage depth in a log-normal distribution channel is essential.
  2. 2.
    The higher the encoding redundancy, the smaller the minimum sequencing coverage depth in any channels.
  3. 3.
    In a log-normal distribution channel, the minimum sequencing coverage depth increases with the number of PCR cycles. This is attributed to varying amplification efficiencies among the designed strands, leading to the over-amplification and sequencing of a small subset of strands as the cycle count rises.
  4. 4.
    The variance reaches its maximum between $\alpha=1$ and $\alpha=2$ for PCR cycles between 10 and 60, and tends to minimize when $\alpha>7$ and $\alpha>8$ respectively, which means that when the expected coverage depth is between 1 and 2, although sequencing $\alpha m$ reads is sufficient to decode all the data, the probability of successfully decoding all the data in a single experiment is minimized. We recommend increasing the sequencing ratio or decreasing the encoding redundancy to reduce the variance.
Figure 3: The expectation of the the coverage depth with respect to $[n,m]$ MDS code redundancy.
Figure 4: The derivative of the variance of the coverage depth. The small figure shows the variance of decoding each oligo under different PCR amplification cycles. The trend of the variance derivative reflects the probability of successfully decoding all the data information in a single sequencing experiment when the expected sequencing sample size is used as the actual sequencing sample size.
IV-B Sequencing Coverage Depth in the Noisy Channel
Theorem 3 and Theorem 4 of this section give two lower bounds on $I_{a}^{\boldsymbol{p}_{t}}(n,m)$ in noisy channels obeying a log-normal distribution.
Remark. In a noisy channel, we assume that at least $a$ noisy copies of each designed strand need to be sequenced for successful retrieval of that strand, i.e., $a>1$.
According to section III, if given the synthetic amount and the amplification efficiency of each strand, denoted as $\boldsymbol{c}=(c_{1},c_{2},\dots,c_{n})$, $\boldsymbol{r}=(r_{1},r_{2},\dots,r_{n})$ respectively, we can deduce the parameters $\mu^{(t)}$ and $\sigma^{(t)}$ of the real log-normal distribution channel after $t$ cycles of PCR.
From [14], we know that if $p_{i}^{(t)}\sim\mathcal{LN}\left(\mu^{(t)},{\sigma^{(t)}}^{2}\right)$, then the expected value of $p_{i}^{(t)}$ is
$\mathbb{E}\left[p_{i}^{(t)}\right]=e^{\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}},$
and the expected value of $\frac{1}{p_{i}^{(t)}}$ is
$\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]=e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}.$
Theorem 3. For given $\boldsymbol{c}$, $\boldsymbol{r}$, for any $t\geq 1$, $\beta>1$, $a>1$, let $R=\frac{n}{m}$, it holds that $P\left[I_{a}^{\boldsymbol{p}_{t}}(n,m)\leq K_{1}\left(\mu^{(t)},\sigma^{(t)},a,R\right)\right]\leq e^{-\beta}\left(1+\frac{m}{n-m}\right)$, where
$K_{1}(\mu^{(t)},\sigma^{(t)},a,R)\triangleq e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\left(\log\Bigg{(}\frac{R}{R-1}\right)-\beta\Bigg{)}.$
Proof. Since $I_{1}^{\boldsymbol{p}_{t}}(n,m)<I_{a}^{\boldsymbol{p}_{t}}(n,m)$, the following proves that
$P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)\leq K_{1}\right]\leq e^{-\beta}\left(1+\frac{m}{n-m}\right).$
Inspired by [7], we have that,
$e^{\log\left(\frac{n}{n-m}\right)-\beta}\cdot\mathbb{E}\left[e^{-\mathbb{E}\left[p_{i}^{(t)}\right]\cdot I_{1}^{\boldsymbol{p}_{t}}(n,m)}\right]$ $=e^{\log\left(\frac{n}{n-m}\right)-\beta}\cdot\sum_{k=1}^{\infty}e^{-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$=\sum_{k=1}^{\infty}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$=\sum_{k=1}^{\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$+\sum_{k=1+\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}^{\infty}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$\geq\sum_{k=1}^{\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$\geq\sum_{k=1}^{\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}1\cdot P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$\geq P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)\leq\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right]$
$=P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)\leq e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\cdot\log\left(\frac{n}{n-m}\right)-e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\cdot\beta\right].$
From [16], the generating function of the geometric random variable $I_{1}^{\boldsymbol{p}_{t}}(n,m)$ is given by
$G_{I_{1}^{\boldsymbol{p}_{t}}(n,m)}\left(e^{-\mathbb{E}\left[p_{i}^{(t)}\right]}\right)$ $=\mathbb{E}\left[{e^{-\mathbb{E}\left[p_{i}^{(t)}\right]\cdot{I_{1}^{\boldsymbol{p}_{t}}(n,m)}}}\right]$
$=\sum_{k=0}^{\infty}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]{\left(e^{-\mathbb{E}\left[p_{i}^{(t)}\right]}\right)}^{k}$
$=\prod_{j=1}^{m}\frac{1-\frac{j-1}{n}}{e^{\mathbb{E}\left[p_{i}^{(t)}\right]}-\frac{j-1}{n}}$
$\leq\prod_{j=1}^{m}\frac{1-\frac{j-1}{n}}{1+\mathbb{E}\left[p_{i}^{(t)}\right]-\frac{j-1}{n}}\leq 1,$
where the inequality is obtained from Taylor expansion of $e^{\mathbb{E}\left[p_{i}^{(t)}\right]}$. Finally, we conclude that,
$P\left[I_{a}^{\boldsymbol{p}_{t}}(n,m)\leq K_{1}\left(\mu^{(t)},\sigma^{(t)},a,R\right)\right]$ $\leq e^{\log\left(\frac{n}{n-m}\right)-\beta}\cdot\mathbb{E}\left[e^{-\mathbb{E}\left[p_{i}^{(t)}\right]\cdot I_{1}^{\boldsymbol{p}_{t}}(n,m)}\right]$
$\leq e^{-\beta}\cdot\frac{n}{n-m}\cdot 1\leq e^{-\beta}\left(1+\frac{m}{n-m}\right).$
[…]
Theorem 3 provides a lower bound on the expected sequencing coverage depth, independent of coding redundancy. The following theorem establishes a lower bound on the expected sequencing coverage depth in relation to coding redundancy. Let $N_{K}$ represents the number of urns that contain less than $a$ balls after $K$ rounds. When using $[n,m]$ MDS code, recovering all data means successfully decoding $m$ out of $n$ strands, implying that at most $n-m$ strands cannot be successfully decoded.
Theorem 4. For given $\boldsymbol{c}$, $\boldsymbol{r}$, for any $a>1$, $t\geq 1$, let $R=\frac{n}{m}$, we have $\mathbb{E}[N_{K}]\leq n-m$, if $I_{a}^{\boldsymbol{p}_{t}}(n,m)>K_{2}\left(\mu^{(t)},\sigma^{(t)},a,R\right)$, where
$K_{2}\left(\mu^{(t)},\sigma^{(t)},a,R\right)\triangleq e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\Bigg{(}(a-1)-\log 2\log\left(\frac{R-1}{R}\right)+(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(\frac{R-1}{R}\right)}\Bigg{)}.$
Before proving Theorem 4, we will first present a lemma, which will be used in the proof of Theorem 4.
Lemma 1. For $K>(a-1)\cdot\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]$, we have $\mathbb{E}\left[N_{K}\right]\leq n-m$, if $\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}e^{-\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}}\leq\frac{1}{e}\left(1-\frac{1}{R}\right)^{\frac{\log 2}{a-1}}$.
Proof. We consider this question in the context of the urn problem. According to Claim 3 in [7], the probability of the $i$-th urn containing less than $a$ balls after $K$ rounds is
$q_{i}=\sum_{j=0}^{a-1}\binom{K}{j}{p_{i}}^{j}(1-p_{i})^{K-j}\leq e^{KD\left(\frac{a-1}{K}\Big{|}\Big{|}p_{i}\right)},$
where
$D(a||b)\triangleq a\log_{2}\frac{a}{b}+(1-a)\log_{2}\frac{1-a}{1-b}$
is the Kullback-Leibler divergence [17]. Then
$\mathbb{E}[N_{K}]=\sum_{i=1}^{n}q_{i}\leq\sum_{i=1}^{n}e^{-(a-1)\log_{2}\frac{a-1}{Kp_{i}}-(K-(a-1))\log_{2}\frac{K-(a-1)}{K\left(1-p_{i}\right)}}\leq ne^{-(a-1)\log_{2}\frac{a-1}{K\mathbb{E}\left[p_{i}\right]}-(K-(a-1))\log_{2}\frac{K-(a-1)}{K\left(1-\mathbb{E}\left[p_{i}\right]\right)}},$
where the second inequality follows from Jensen’s inequality.
Hence, $\mathbb{E}[N_{K}]\leq n-k=n(1-R)$ holds if and only if
$ne^{-(a-1)\log_{2}\frac{a-1}{K\mathbb{E}\left[p_{i}\right]}-(K-(a-1))\log_{2}\frac{K-(a-1)}{K\left(1-\mathbb{E}\left[p_{i}\right]\right)}}\leq n(1-R),$
that is,
$e^{-(a-1)\log_{2}\frac{a-1}{K\mathbb{E}\left[p_{i}\right]}-(K-(a-1))\log_{2}\frac{K-(a-1)}{K\left(1-\mathbb{E}\left[p_{i}\right]\right)}}\leq 1-R.$ (6)
By the change of base formula, the following equality holds.
$e^{-(a-1)\log_{2}\frac{a-1}{K\mathbb{E}\left[p_{i}\right]}-(K-(a-1))\log_{2}\frac{K-(a-1)}{K\left(1-\mathbb{E}\left[p_{i}\right]\right)}}=\left(\frac{K\mathbb{E}\left[p_{i}\right]}{a-1}\right)^{\frac{a-1}{\ln{2}}}\left(\frac{K\cdot\frac{1}{\mathbb{E}\left[p_{i}\right]}-K}{K\cdot\frac{1}{\mathbb{E}\left[p_{i}\right]}-(a-1)\cdot\frac{1}{\mathbb{E}\left[p_{i}\right]}}\right)^{\frac{K-(a-1)}{\ln{2}}}.$
Let $K\triangleq\lambda(a-1)\cdot\frac{1}{\mathbb{E}\left[p_{i}\right]}$, then (6) can be rewrite as
$\lambda\left(1-\frac{\lambda-1}{\lambda\cdot\frac{1}{\mathbb{E}\left[p_{i}\right]}-1}\right)^{\lambda\cdot\frac{1}{\mathbb{E}\left[p_{i}\right]}-1}\leq(1-R)^{\frac{\ln{2}}{a-1}}.$ (7)
According to the definition of e, we know that
$\left(1-\frac{\lambda-1}{\lambda\cdot\frac{1}{\mathbb{E}\left[p_{i}\right]}-1}\right)^{\lambda\cdot\frac{1}{\mathbb{E}\left[p_{i}\right]}-1}\leq e^{-(\lambda-1)}.$
Thus, if $\lambda e^{-\lambda}\leq\frac{1}{e}(1-R)^{\frac{\ln{2}}{a-1}}$ holds then (7) also holds.
By the assumption,
$\frac{K\mathbb{E}\left[p_{i}\right]}{a-1}e^{\frac{K\mathbb{E}\left[p_{i}\right]}{a-1}}\leq\frac{1}{e}(1-R)^{\frac{\ln{2}}{a-1}}.$
Based on Lemma 1, we will prove Theorem 4 next.
Proof of Theorem 4. Let $x=\frac{1}{e}\left(1-\frac{1}{R}\right)^{\frac{\log 2}{a-1}}$, $y=\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}$, according to Claim 7 in [7], the equation
$-\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}e^{-\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}}=-ye^{-y}=-x=-\frac{1}{e}\left(1-\frac{1}{R}\right)^{\frac{\log 2}{a-1}}$
has two solutions $y=-W_{0}(-x)$ and $y=-W_{-1}(-x)$, and Lemma 1 holds if and only if $y\geq-W_{-1}(-x)$. Then, from Claim 8 in [7], we have that,
$-W_{-1}(-x)<1+\sqrt{-\frac{2\log 2}{a-1}\log\left(1-\frac{1}{R}\right)}-\frac{\log 2}{a-1}\log\left(1-\frac{1}{R}\right).$
Hence, according to the expression of $y$, a sufficient condition for $\mathbb{E}\left[N_{K}\right]\leq n-m$ is
$\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}\geq 1+\sqrt{-\frac{2\log 2}{a-1}\log\left(1-\frac{1}{R}\right)}-\frac{\log 2}{a-1}\log\left(1-\frac{1}{R}\right),$
equals to
$K\geq\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\cdot(a-1)-\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\cdot\log 2\log\left(1-\frac{1}{R}\right)+\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\cdot(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(1-\frac{1}{R}\right)}.$
According to Jensen’s inequality, $\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\leq\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]$, thus a more sufficient condition for $\mathbb{E}\left[N_{K}\right]\leq n-m$ is
$K\geq\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot(a-1)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log 2\log\left(1-\frac{1}{R}\right)+\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(1-\frac{1}{R}\right)},$
that is,
$e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\left((a-1)-\log 2\log\left(\frac{R-1}{R}\right)+(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(\frac{R-1}{R}\right)}\right).$
[…]
We will visually compare the two lower bounds of the expected sequencing coverage depth that we prove with the maximum sequencing coverage depth in the real channel obtained through 1000 Monte Carlo simulations, as shown in the fig. 5 below.
(a) R=1.25 a=2
(b) R=2 a=2
Figure 5: Visualization of the theoretical lower bounds of the expected sequencing coverage depth and 1000 Monte Carlo simulations of the maximum sequencing coverage depth in the real channel.
V Conclusion
In this paper, by working with PCR and sequencing data, we analyze and simulate the probability distribution of the real channel, and mainly investigate the sequencing coverage depth problem based on experimental data, under the non-random access setting. That is, we prove the expected coverage depth and its theoretical lower bounds in real noiseless and noisy channels respectively, and first propose the problem of decoding all data successfully in a single sequencing experiment using the expected coverage depth as the sample size.